### v1

# revealing an in-context neural scaling law

Abstract
Pretrained large language models (LLMs) are surprisingly effective at performing zero-shot tasks, including time-series forecasting. However, understanding the mechanisms behind such capabilities remains highly challenging due to the complexity of the models. In this paper, we study LLMs’ ability to extrapolate the behavior of dynamical systems whose evolution is governed by principles of physical interest. Our results show that LLaMA 2, a language model trained primarily on texts, achieves accurate predictions of dynamical system time series without fine-tuning or prompt engineering. Moreover, the accuracy of the learned physical rules increases with the length of the input context window, revealing an in-context version of neural scaling law. Along the way, we present a flexible and efficient algorithm for extracting probability density functions of multi-digit numbers directly from LLMs.
1 Introduction
Large language models (LLMs) are billion-parameter neural networks that have been trained on a vast corpus of text to perform next-word predictions (Radford et al., 2019; Brown et al., 2020; Kaplan et al., 2020; OpenAI, 2023; Touvron et al., 2023). Since the introduction of the transformer architecture (Vaswani et al., 2017), LLMs have become remarkably proficient in a wide range of natural language processing (NLP) tasks, including question answering (Roberts et al., 2020; Drori et al., 2022) and translation (Brown et al., 2020). In the process of training larger models, unexpected emergent properties of LLMs have been discovered (Wei et al., 2022a), such as program execution (Nye et al., 2021) and multi-step reasoning (Cobbe et al., 2021; Wei et al., 2022b; Suzgun et al., 2022).
While powerful, LLMs remain scientifically mysterious. How they compress vast amounts of information and implement complex algorithms within their architecture is not readily discoverable. To attempt to shine some light on these “black box” models, recent studies notably used representation probes to decipher some of the mechanisms encoded in the layers of trained neural networks (Alain & Bengio, 2016; Akyürek et al., 2022; Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Nanda et al., 2023; Park et al., 2023). By showing, experimentally, that features of prompt representations in a transformer’s latent space can be mapped onto human-comprehensible concepts (e.g., statement veracity, geographical localization, etc.), these findings suggest the possibility of an internal “world model”, which provides LLMs with some sort of perception of reality, far more sophisticated than that of a mere “stochastic parrot.”
In this work, we continue these explorations of intrinsic world models by proposing a new perspective and empirical approach. Specifically, we are interested in the (apparent) unreasonable effectiveness of LLMs in understanding the physical world. We then leverage and extend recent observations that LLMs are capable of time series extrapolation without specific prompting or fine-tuning (Gruver et al., 2023; Jin et al., 2023a). Instead of forecasting empirical time series, we study the ability of LLMs to extrapolate dynamical systems time series while maintaining fidelity with underlying, governing principles.
Figure 1: Continuing a non-linear pendulum angular trajectory using LLaMA-13b. The prediction accuracy increases with the input context length. The temperature parameter is set to $10^{-7}$ to eliminate randomness in auto-regressive continuation.
As an example, Figure 1 illustrates an LLM’s ability to learn in-context and continue the deterministic trajectory of a non-linear pendulum, whose amplitude and frequency both drift over time (Section A.4). The prediction accuracy improves with the length of the input time series, which we refer to as context length.
We focus on systems that are either deterministic or stochastic but have well-defined transition rules between time steps. We compare the transition rules inferred by the LLM with the ground truth underlying the input data by using specific loss functions. We find that loss functions decrease rapidly with the number of states observed in-context, revealing a neural scaling law for in-context learning, as illustrated in Figure 2. This observed behavior spans a wide range of dynamical systems, from physically-grounded systems like the Lorenz system to entirely synthetic ones, such as Markov chains with arbitrary transition probabilities. Our results suggest that LLMs might possess not only an internal model of the world (Gurnee & Tegmark, 2023), but moreover an in-context learning ability which seems to allow them to glean physical insights from input data.
Figure 2: Evolution of the loss function for the predicted next state by LLaMA-13b with respect to the number of observed states in various physical systems. We employ the Bhattacharyya distance as a loss function for stochastic systems (solid lines), and the squared deviations from the mean (SDM) for deterministic systems (dashed lines). Brownian motion and geometric Brownian motion deviate significantly from power law scaling, which can be explained by their lack of stationary distributions (Section A.2).
Organization of the paper. We begin in Section 2 by providing background information on LLMs for time series forecasting. We then describe our methods for tokenizing time series and extracting probability density functions from LLMs in Section 3. Our experiments and their results on various dynamical systems using the LLaMA model (Touvron et al., 2023) are presented in Section 4. We conclude in Section 5 by discussing the implications of our findings for the understanding of in-context learning abilities and the intrinsic world model of LLMs.
Main contributions. The main contributions of this paper are summarized as follows:
  • demonstrating LLMs’ zero-shot ability to model the evolution of dynamical systems without instruction prompting;
  • implementing a computationally efficient framework called Hierarchy-PDF to extract statistical information of a dynamical system learned by a transformer-based LLM;
  • discovering a scaling law between the accuracy of the learned transition rules and the context window length.
2 Background and related work
LLMs are primarily designed, trained, and utilized for NLP, i.e. the ability to generate meaningful language in response to a textual prompt. However, scaled-up modern models have been found to perform surprisingly well for tasks beyond NLP (Wei et al., 2022a).
Among these emergent capabilities, a specific class called in-context learning, consists of learning from and extrapolating sequences included in the prompt, even when they are not language (Brown et al., 2020). One useful example of in-context learning is zero-shot time series forecasting (Gruver et al., 2023). Given a numerical time series of sufficient length, LLMs can extrapolate the prompt sequence and generate numerical values that are consistent with the input without any fine-tuning, and sometimes more accurately than specialized algorithms for time series extrapolation (Wu et al., 2021; Zhou et al., 2022).
The work of (Gruver et al., 2023) aims to forecast empirical time series and introduces a tokenization procedure to convert a sequence of floating point numbers into appropriate textual prompts for LLMs. This led to several subsequent studies on the application of LLMs for time series forecasting (Chen et al., 2023; Jin et al., 2023b, a; Dooley et al., 2023; Schoenegger & Park, 2023; Wang et al., 2023; Xu et al., 2023).
Unlike these prior studies, our paper does not focus on forecasting real-world time series, such as weather data (Zhou et al., 2021) or electricity demand (Godahewa et al., 2021), where the underlying model generating the sequence is unavailable or undefined. Instead, we aim to extract the learned transition rules from the probability vector generated by the LLM and compare them against the ground truth rules (chaotic, stochastic, discrete, continuous, etc.) governing the input time series.
3 Methodology
Our methodology for testing LLMs’ ability to learn physical rules from in-context data consists of three steps:
  1. 1.
    Sample a time series $\{x_{t}\}_{t\geq 0}$ from a given dynamical system governed by Markovian transition rules $P_{ij}$.
  2. 2.
    Prompt the LLM with this time series to extract the learned probability densities for subsequent digits $\tilde{P}_{ij}$.
  3. 3.
    Measure the discrepancy between the ground truth $P_{ij}$ and learned $\tilde{P}_{ij}$ using appropriate loss functions.
3.1 Prompt generation
Markov processes. Most of our testing data may be modeled as discrete-time Markov chains, where the PDF of the next state at time $t+1$ depends solely on the previous state $x_{t}$ at time $t$:
$P(X_{t+1}|X_{1}=x_{1},\dots,X_{t}=x_{t})=P(X_{t+1}|X_{t}=x_{t}).$
This models either discrete iterative systems or continuous dynamical systems after time-discretization using a first-order explicit time-stepping scheme. Each time series is governed by a set of transition rules $P(X_{t+1}|X_{t},\Delta t)$, which, in general, depends on both the previous state $x_{t}$ and the time step $\Delta t$. However, we assume a uniform time-discretization of the system (i.e., $\Delta t$ is constant), and consider time-independent transition rules: $P(X_{t+1}|X_{t})$. Hence, we are interested in how LLMs learn transition rules in-context, as opposed to forecasting unseen transition rules. It is worth noting that the Markovian nature does not imply that the transformer only relies on $x_{t}$ to predict $X_{t+1}$. Instead, an efficient model might combine all previous states $\{x_{0},\ldots,x_{t}\}$ to hypothesize a reasonable transition rule $P(X_{t+1}|X_{t})$, which can then be used, in conjunction with the last observed state $x_{t}$ to formalize a probabilistic prediction of $X_{t+1}$.
Time series tokenization. An input time series typically consists of ($\sim 10^{3}$) time steps, each represented as a real number. We first rescale each number and represent it using $n$ digits (typically, $n=3$). Each time series is rescaled to the interval $[1.50,8.50]$ so that the number of digits never changes throughout the series. In particular, there is an intentional gap between the upper limit $8.5$ and $10$, so that if an LLM outputs $1$ as the first digit, it is to be unambiguously interpreted as a number $1.xx$, as opposed to $1x.x$. Similarly, the lower limit of $1.5$ further eliminates ambiguity by circumventing leading zeros.
We then follow the scheme introduced in (Gruver et al., 2023; Jin et al., 2023a) to serialize the time series as strings and tokenize them. Since our numerical experiments are performed with LLaMA models (Touvron et al., 2023) (though our conclusions apply to other models, see Appendix A), which contain ten distinct tokens for the digits $0,\ldots,9$, we pass the string-digits without spacing but with a comma separating time steps. In the top right of Figure 4, we show an example of a time series serialized as an input string.
3.2 Extraction of transition rules
Discrete state space. When the Markov process is discrete and has a finite state space, each state can be represented by a single token. While one could theoretically select any subset within the set of 32k tokens in the LLaMA codebook, we employ tokens corresponding to the ten number strings: $0,\ldots,9$. Hence, these strings are represented with ten independent tokens, while letter tokens are often combined to form sub-components of words. Moreover, we observe in Section 4 that the most sophisticated LLaMA model (LLaMA-70b) that we evaluate can only learn up to 9 discrete states. Therefore, we do not attempt to go beyond 9 distinct states by extending to non-number tokens.
Figure 3: Extracting learned transition rules of systems with discrete state space.
Figure 3 illustrates our framework for learning discrete Markov chains with LLMs. First, we randomly draw an $n\times n$ matrix $(P_{ij})$ from the space of transition matrices. We then sample the Markov chain, tokenize the series and prompt the LLM with it. The length of the series is chosen such that the tokenized representation does not exceed the length of the LLM’s context window. We extract the LLM’s prediction for the next state by performing a softmax operation on the output logits corresponding to the $n$ allowed states and discarding all other logits. This is equivalent to modifying the non-state logits using an additive bias of negative infinity, which is consistent with the sampling method described in (Gruver et al., 2023; Jin et al., 2023a).
Continuous-time state space. Stochastic processes, such as the Brownian motion (Einstein, 1905; Perrin, 1909), might have continuous-time state space $\{X_{t}\}_{t\geq 0}$, indexed by a real variable $t\geq 0$. For these processes, we represent the value of each state as a multi-digit number and separate each state using the comma symbol “,”, which also has its own token in the LLaMA codebook, and is never lumped along with neighboring numbers. As observed in (Jin et al., 2023a), an LLM prediction of multi-digit values can be naturally interpreted as a hierarchical softmax distribution (Mnih & Hinton, 2008; Challu et al., 2023).
Figure 4: Extracting transition rules of systems with continuous state space. The learned transition rules are more refined for bins near the mode, which dominate the Bhattacharyya distance. The height of $\tilde{P}(x_{t+1}|x_{t})$ is shown in log scale.
Specifically, let $u$ denote a multi-digit string representing the value of a state at a given time-step, then the LLM’s softmax prediction for the $i^{th}$ digit, $u_{i}$, provides a histogram of ten bins of width $0.1^{i}$. Subsequently, the prediction of the $(i+1)^{\mathrm{th}}$ digit goes down one level into the hierarchical tree by refining one of the bins into ten finer bins of width $0.1^{i+1}$, and so on until the last digit is processed (see Figure 4).
Hierarchy-PDF algorithm. While a single pass through the LLM yields a discretized PDF represented by bins of various widths, we can refine the PDF by querying each coarse bin. For example, to obtain a maximal resolution PDF of a 3-digit value, we need to query all $10^{2}$ combinations of the first two digits and furnish a PDF of uniform bin width, $0.1^{3}$. Suppose a time series consists of S values (steps), each represented as $n$ digits, then furnishing maximal precision PDFs for the entire sequence requires $10^{n-1}S$ forward passes of the LLM. This seemingly daunting complexity is significantly reduced because most of the $10^{n-1}S$ inputs differ only in the last tokens, and thus one can recursively cache the key and value matrices associated with the shared tokens. The complexity is further reduced by refining only the high-probability bins near the mode, as shown in Figure 4. Algorithm 1 outlines the Hierarchy-PDF algorithm used to recursively refine the PDF associated with a multi-digit value in a time series. More algorithmic details are available in Appendix A as well as in the associated GitHub repository.
Algorithm 1 Hierarchy-PDF
  Input: Unrefined PDF, current depth $D_{c}$, target depth $D_{t}$
  Procedure: RecursiveRefiner($D_{c}$, $D_{t}$)
     if $D_{c}=D_{t}$ then
        end the recursion
     else if current branch is refined then
        Change the last digit to launch 9 recursive branches
        RecursiveRefiner($D_{c}$, $D_{t}$)
     else if current branch is unrefined then
        refine multi_PDF with new logits
        if $D_{c}+1<D_{t}$ then
           Append digits to launch 10 recursive branches
           RecursiveRefiner($D_{c}+1$, $D_{t}$)
        end if
     end if
3.3 Loss functions
Once the learned transition rules, $\tilde{P}(X_{t+1}|X_{t})$, have been extracted, we quantify the deviation from the ground truth $P(X_{t+1}|X_{t})$. Depending on the nature of the system, one of the following two loss functions may be more appropriate (see Section 4).
Bhattacharyya distance. For stochastic time series, we use the Bhattacharyya distance to characterize the distance between learned and ground truth transition functions. The Bhattacharyya distance between $P$ and $\tilde{P}$, on a domain $\mathcal{X}$ is defined as (Bhattacharyya, 1943, 1946; Kailath, 1967):
$\mathrm{D_{B}}(P,\tilde{P})=-\ln\left(\int_{\mathcal{X}}\sqrt{p(x)\tilde{p}(x)}\,\textup{d}x\right),$ (1)
and has been widely employed by feature selection and signal extraction methods (Choi & Lee, 2003; Kailath, 1967). Since $\tilde{P}(X_{t+1}|X_{t})$ takes the form of a hierarchical PDF, one may approximate the integral in Equation 1 via a discrete quadrature rule as
$\mathrm{D_{B}}(P,\tilde{P})=-\ln\left(\sum_{x}\sqrt{p(x)\tilde{p}(x)}\Delta x\right),$ (2)
where $\Delta x$ denotes the length of the sub-interval containing $x$ in the partition of $\mathcal{X}$.
Squared deviations from the mean. For deterministic systems, the true transition functions become delta-functions. As a result, the discretized Bhattacharyya distance from Equation 2 reduces to (see Equation 18 in Section A.6)
$\mathrm{D_{B}}(\delta(x-x_{\mathrm{true}}),\tilde{P})=-\frac{1}{2}\ln(\tilde{p}(x_{\mathrm{true}}))+C,$
which is proportional to the negative log-likelihood (NLL) assigned to the true data by the LLM, plus a constant $C$ [^1] This constant is determined by the base B of the system, and the number of digits n as $C=-\ln\Delta x=n\log B$.. NLL references only the finest bins in the hierarchical PDF and is thus unstable as in-context loss. As an alternative, we use the squared deviations from the mean (SDM) (Kobayashi & Salam, 2000) as the in-context loss for deterministic systems:
$\mathrm{SDM}(x_{\mathrm{true}},\tilde{P})=\left(x_{\mathrm{true}}-\sum_{x\in\mathcal{X}}\tilde{p}(x)x\Delta x\right)^{2},$
where the mean $\mu_{\tilde{P}}=\sum_{x}\tilde{p}(x)x\Delta x$ is extracted from the hierarchical PDF. Note that unlike the Bhattacharyya distance, which references the model prediction $\tilde{p}$ only at $x_{\mathrm{true}}$, the SDM takes into account the entire support $x\in\mathcal{X}$. Our numerical experiments suggest that SDM is more stable and better captures the in-context learning dynamics of deterministic systems (see Section 4.1).
4 Experiments and Results
This section reports empirical in-context learning results on several dynamical systems modeling physics phenomena, such as a Brownian motion, the logistic map with chaotic behavior, and the Lorenz system. All experiments are repeated with ten trajectories initiated by different random seeds for stochastic processes, or initial conditions for chaotic systems. In each case, we extract the transition rules $P(X_{t}|X_{t-1})$ learned by the model, and then calculate the in-context loss curves, using the Hierarchy-PDF algorithm described in Section 3.
Choice of the model. All numerical experiments reported in this section are performed using the LLaMA-13b model (Touvron et al., 2023), which is open source, unlike GPT-4, and does not have the tokenization issue associated with GPT-3 (Gruver et al., 2023). While we observe that larger language models such as LLaMA-70b may achieve lower in-context loss on some dynamical systems (Section A.5), they do not display qualitative differences.
4.1 Deterministic Systems
4.1.1 Logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behavior (Strogatz, 2015). It is governed by the following iterative equation:
$x_{t+1}=f(x_{t})=rx_{t}(1-x_{t}),\quad x_{0}\in(0,1),$ (3)
which may also be written using conditional distributions to reflect the deterministic nature of the system as
$X_{t+1}|\{X_{t}=x_{t}\}\sim\delta_{f(x_{t})},$
where $\delta$ denotes the Dirac delta distribution. This conditional distribution is the ground truth transition function displayed in red in Figure 5. The parameter $r\in[1,4)$ controls the behavior of the system and is set to $r=3.9$. At this value, the dynamics are naturally confined within the interval $(0,1)$, and the system has no stable fixed points. Due to the chaotic nature of the system, two initial nearby trajectories diverge exponentially in time. This property allows us to sample multiple uncorrelated trajectories by using different initial conditions $x_{0}$, sampled uniformly in $(0,1)$.
Figure 5: Next state prediction of the logistic map. Top: Input chaotic time series shown in black, and the state to be predicted is highlighted in red. Bottom: The LLM’s statistical prediction for the last state. The ground truth distribution is delta-distributed, which is shown as a vertical red line.
Figure 5 displays one of the ten tested trajectories and an LLM’s prediction of the last state. The PDF of the next state prediction is extracted using the Hierarchy-PDF algorithm described in Section 3.2. We find that the LLM prediction is close to the ground truth, except for minor deviations manifested by small, but non-zero, probability densities in neighboring values. While the extracted prediction is only reported for the last time step in the bottom panel of Figure 5, we also extract the model prediction at every time step for all tested trajectories and report the corresponding Bhattacharyya and SDM losses in Figure 6.
Figure 6: Logistic map in-context loss curves. For deterministic systems, Bhattacharyya loss is subject to large fluctuations while SDM loss is more stable.
As foreshadowed in Section 3, the Bhattacharyya loss suffers from large fluctuations with deterministic systems such as the logistic map, while the SDM loss better captures the in-context learning dynamics. In particular, the SDM loss decreases rapidly with the number of observed states, without any fine-tuning nor prompt engineering of the LLM. This suggests that the LLM can extract the underlying transition rules of the logistic map from in-context data.
4.1.2 Lorenz system
The Lorenz system (Lorenz, 1963) is a three-dimensional (3D) dynamical system derived from a simplified model of convection rolls in the atmosphere. It consists of a system of three ordinary differential equations:
$\dot{x}(t)$ $=\sigma(y-x),\quad\dot{y}(t)=x(\rho-z)-y,$
$\dot{z}(t)$ $=xy-\beta z,$
where $\sigma=10$, $\rho=28$ and $\beta=8/3$ are parameters dictating the chaotic behavior of the system. We compute ten 3D trajectories using a first-order explicit time-stepping scheme. All trajectories share the same initial conditions in $y$ and $z$, and differ only in the $x$-coordinate, which is uniformly sampled in $(0,0.3)$. The chaotic nature of the system guarantees that the sampled trajectories quickly diverge from one another. We prompt the LLM with the $x$-component of the simulated series and extract the next predicting values.
Figure 7: Loss curves for predicting the $x$-component of the Lorenz system with respect to the number of observed time steps.
When the $x$, $y$, and $z$ components are observed, the system is deterministic and Markovian; in the sense that a state vector $\vec{s}_{t}=(x_{t},y_{t},z_{t})$ at time $t$ fully determines the next state $\vec{s}_{t+1}$. However, if the $x$-component is the only one observed, then the system ceases to be Markovian but remains deterministic if one expands the state vector to include information from earlier states. Hence, Takens’ embedding theorem (Takens, 2006) guarantees that the observation of at most seven states of the series $x_{0:t}$ is sufficient to predict $x_{t+1}$. Finding the optimal number of states to reconstruct the system’s trajectory is an area of active research (Strogatz, 2015). Despite this apparent difficulty, LLaMA-13b can formulate increasingly accurate predictions of the series as it observes more states, as evidenced by the decaying loss curves plotted in Figure 7.
4.2 Stochastic Systems
4.2.1 Brownian motion
Brownian motion is an example of a continuous-time stochastic process (Einstein, 1905), and is described by a stochastic differential equation (SDE):
$dX_{t}=\mu dt+\sigma dW_{t},$ (4)
where $X_{t}$ represents the state of the system at time $t$, $\mu$ is the drift coefficient, $\sigma$ is the volatility coefficient, and $dW_{t}$ is the increments of a Wiener process (Revuz & Yor, 2013), modeling the randomness of motion.
Figure 8: Next state prediction of Brownian motion. Top: Input stochastic time series shown in black, and the state to be predicted is highlighted in red. Bottom: The LLM’s prediction, along with the ground truth distribution.
To simulate trajectories of Brownian motion, we use the Euler–Maruyama method (Platen, 1999), which discretizes Equation 4 as $X_{t+\Delta t}=X_{t}+\mu\Delta t+\sigma\sqrt{\Delta t}Z$, where $\Delta t$ is the time resolution, and $Z\sim\mathcal{N}(0,1)$ is a random variable that follows a standard Gaussian distribution. The Euler–Maruyama method may also be written as a conditional distribution:
$X_{t+\Delta t}|\{X_{t}=x_{t}\}\sim\mathcal{N}(x_{t}+\mu\Delta t,\sigma^{2}\Delta t),$
which is the ground truth transition function visualized in Figure 8. Indeed, the ground truth next state is described as a Gaussian distribution, and we observe in Figure 8 that the LLM prediction agrees well with the true, underlying distribution. Additionally, as shown in Figure 8, the LLM displays the correct Gaussian shape for the PDF, converging to a measured kurtosis of $3$ (see Section A.6). We then simulate ten different trajectories using random seeds for $Z$ and report the resulting LLM learning curves in Figure 9, measured in the Bhattacharyya distance.
Figure 9: Bhattacharyya distance between the LLM predicted PDF and the ground truth transition function of Brownian motion with respect to the input length.
4.2.2 Geometric Brownian motion
Geometric Brownian motion (GBM) (Oksendal, 2013) is a stochastic process that is commonly used in mathematical finance to model the trajectories of stock prices and other financial assets (Hull, 2021). A GBM is governed by the following SDE:
$dX_{t}=\mu X_{t}dt+\sigma X_{t}dW_{t},$ (5)
where $X_{t}$ models the price of an asset at time $t$, and the fluctuation term $\sigma X_{t}dW_{t}$ is proportional to the current asset price $X_{t}$. The Euler–Maruyama discretization of the GBM reads $X_{t+\Delta t}=X_{t}+\mu X_{t}\Delta t+\sigma X_{t}\sqrt{\Delta t}Z$, and leads to the ground truth transition function:
$X_{t+\Delta t}|\{X_{t}=x_{t}\}\sim\mathcal{N}(x_{t}+\mu x_{t}\Delta t,(\sigma x_{t})^{2}\Delta t).$ (6)
We simulate ten different GBM trajectories using random seeds and report the corresponding learning curves in Figure 10.
Figure 10: Geometric Brownian motion in-context loss curve.
We perform an additional numerical test to verify that the LLM is learning the correct relationship between the variance of the GBM and the state value $X_{t}$ (see Equation 6). To investigate this, we display in Figure 11 the expected standard deviation along with the learned one, extracted from the Hierarchy-PDF using Equation 20, across all predicted states. We find that the LLM respects the ground truth standard deviation of the GBM, as prescribed by the underlying transition function.
Figure 11: Evolution of the geometric Brownian motion standard deviation with respect to the state value $x_{t}$ (see Equation 6), along with the predicted standard deviation extracted from the LLM at each time step.
4.2.3 Noisy logistic map
The logistic map system becomes stochastic when one introduces small Gaussian perturbations of variance $\sigma^{2}$ at each step to model noisy measurements: $x_{t+1}=f(x_{t}+\epsilon)$ where $\epsilon\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. This implies that the conditional distribution has non-zero uncertainty and is no longer delta-like [^2] The approximation in Equation 7 assumes a small perturbation compared to the second derivative, that is $\sigma^{2}\ll 1/f^{\prime\prime}(x)$.:
$X_{t+1}|\{X_{t}=x_{t}\}\sim\mathcal{N}\left(f(x_{t}),(\sigma f^{\prime}(x_{t}))^{2}\right).$ (7)
The first derivative of $f$ measures how sensitive the local dynamics are to external perturbations. This intuitively explains why the standard deviation of the conditional distribution should be proportional to $f^{\prime}$.
Figure 12: Stochastic logistic map in-context loss curves.
Figure 13: Noisy logistic map standard deviation as a function of the state value $x_{t}$, learned by the LLM, along with the ground truth.
We observe a fast decay of the in-context loss function with respect to the length of the observed time series in Figure 12. This is in agreement with previous numerical experiments and suggests that the LLM’s prediction ability is robust against noise perturbation. Moreover, the correct model for the standard deviation of the time series is encoded in the LLM PDF, as shown in Figure 13, and can be extracted with the Hierarchy-PDF algorithm.
4.2.4 Markov chains with discrete states
The transition rules of a time-independent Markov chain with $n$ states consist of a stochastic matrix $(P_{ij})_{1\leq i,j\leq n}$, defined as
$P_{ij}=P(X_{t+1}=j|X_{t}=i),\quad 1\leq i,j\leq n.$
Using the testing procedures elaborated in Section 3.2, we generate ten Markov chains, each from a distinct and randomly generated transition matrix of size $n=4$.
Figure 14: Markov chain in-context loss curves decay rapidly with respect to the input time series length.
The corresponding loss curves between the LLM predictions and the ground truth are displayed in Figure 14. Unlike the previous dynamical systems investigated earlier, which are all motivated by models pertaining to physical phenomena, the discrete Markov chains are commonly used to describe the evolution of symbolic systems, such as language (Almutiri & Nadeem, 2022) and decision-making process (Feinberg & Shwartz, 2012). In our experiments, the transition rules are synthesized completely at random. Yet, the LLM formulates remarkably accurate statistical predictions as more data are observed in-context. These conclusions hold for larger transition matrices ($n>4$) and bigger LLMs, such as LLaMA-70b (see Section A.5).
5 Discussion and conclusion
We discuss the limitations, interpretations, and implications of the experimental findings presented in this work.
In-context neural scaling law. Neural scaling laws (Kaplan et al., 2020) are power laws that characterize how the loss of trained neural networks vary with respect to parameters of the model, such as model size, dataset size, and computational resources. To the best of our knowledge, neural scaling laws have, so far, only been observed in the training procedure, which explicitly updates the weights of neural networks using optimization algorithms. The loss curves observed in the different numerical experiments (see Figure 15 in Section A.1) reveal an additional in-context scaling law for LLMs zero-shot learning of dynamical systems.
Early plateauing and stationary distributions. While most of our experimental datasets give rise to power scaling laws, two loss curves —the Brownian motion and geometric Brownian motion— “plateau early” at a context length of approximately $10^{2}$ (see Figure 15). One potential explanation for this phenomena, elaborated in Section A.2, is that the Brownian and geometric Brownian motions “wander out of distribution” at large time $t>0$, while all the other time series associated with the dynamical systems studied in this paper converge to stationary distributions.
Temperature. The temperature hyperparameter $T$ affects the outcome of the LLM softmax layer and, hence, the reconstructed PDF obtained through the Hierarchy-PDF algorithm. The next digits prediction are extracted at the parameter $T=1$, which is the default temperature used during training (Touvron et al., 2023). However, we found that using a temperature of $T=0.7$ might be more suitable for extracting predicted specific statistical parameters such as the standard deviation from the PDF (see Figures 11 and 13, and the discussion in Section A.3).
Data leakage. The possibility that LLMs’ accurate predictions of the next time step value are merely due to memorization seems extremely unlikely. While we cannot rule out the existence of numerical sequences in the training corpus (especially popular ones like the Fibonacci sequence or the digits of $\pi$), our experiments rely on custom time series with arbitrary parameters. In addition, these sequences have been normalized for proper tokenization. A sequence of even one thousand numerical values, encoded with three digits, may cover $10^{3000}$ particular instances, which is well beyond the $\sim 10^{12}$ tokens of the training corpus (Touvron et al., 2023). Therefore, data leakage in the context of numerical time series appears implausibly unlikely.
In conclusion, we showed that, with increasing cardinality in tokenization, and given sufficient context, LLMs can accurately extrapolate not just deterministic trajectories (non-linear pendulum), but also chaotic and stochastic sequences governed by specific transition rules. In the latter case, the extrapolation accuracy is measured in a statistical sense. This suggests that large language models can in-context learn dynamical systems’ time series and predict future states in a manner that maintains fidelity with the underlying principles governing the system’s evolution. Moreover, this behavior is observed without the use of any fine-tuning or “prompting engineering”.
The implications are substantial. Our results suggest that a transformer, trained primarily on textual data, can extract governing principles of numerical sequences observed in-context. This further supports the notion that LLMs may somehow encode in their architecture, a rich and detailed representation of the world. Prior investigations of the “world model” hypothesis (Gurnee & Tegmark, 2023; Li et al., 2023; Patel & Pavlick, 2021) primarily focus on internal representations of static semantic concepts. In contrast, our study indicates that an LLM’s world model also encompasses the mathematical principles underlying the dynamics of both physical systems, like the logistic map, and symbolic systems, such as Markov chains.
An exciting future direction is to study the generalization of the observed in-context neural scaling laws for other LLMs, such as GPT4 (OpenAI, 2023), and the newly proposed state space models (Gu & Dao, 2023). Furthermore, the in-context neural scaling law might hint at a learning algorithm that LLMs implicitly implement during inference (e.g., a numerical time-stepping scheme), and characterizing such an algorithm is an open question of broad interest.
Software and Data
The code and data supporting this study will be deposited on a public GitHub repository upon acceptance.
Broader Impact
While there are several potential societal consequences of large language models, the main goal of this work is to advance the field of machine learning.
Acknowledgements
This work was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Numbers N00014-23-1-2729 and N00014-23-1-2716.
Appendix A Appendix
A.1 In-context neural scaling law
Neural scaling laws (Kaplan et al., 2020) describe how the performance of trained neural networks, particularly language models, scales with changes in key factors such as model size ($N$), dataset size ($D$), and computational resources used for training (C). These laws are often observed as power-law relationships in the following form:
$L(N)$ $=\left(\frac{N}{N_{\text{c}}}\right)^{\alpha_{N}},$ (8)
$L(D)$ $=\left(\frac{D}{D_{\text{c}}}\right)^{\alpha_{D}},$ (9)
$L(C)$ $=\left(\frac{C}{C_{\text{c}}}\right)^{\alpha_{C}},$ (10)
where $L$ represents the loss or performance metric of the model. The characteristic factors ($N_{\text{c}}$, $D_{\text{c}}$, and $C_{\text{c}}$), and power coefficient ($\alpha$) are extracted empirically from training curves. The fitted quantities depend on the distribution of data, the model architecture and the type of optimizer used for training. Such power-law relations appear in log-log plots as straight lines, whose slopes correspond to the parameter $\alpha$.
Our loss curves from learning dynamical systems (see Figure 2) reveal an additional neural scaling law that applies to in-context learning:
$L(D_{\text{in}})=\left(\frac{D_{\text{in}}}{D_{\text{c}}}\right)^{\alpha},$ (11)
where $D_{\text{in}}$ stands for the length of time series observed in the prompt (in-context). In Figure 15, we display the fitted power laws to the in-context loss curves.
Figure 15: In-context loss curves from LLaMA-13b fitted with power law, with fitted power coefficient $\alpha$ shown in legend. Left: loss of stochastic series measured in Bhattacharyya distance. Right: loss of deterministic series measured in SDM.
To the best of our knowledge, neural scaling laws have so far only been observed in the training procedure, which updates the weights of neural networks using an explicit algorithm, such as stochastic gradient descent and Adam (Kingma & Ba, 2017).
A.2 Invariant measure and the early plateauing of in-context loss
While most datasets are well-described by the power laws, two loss curves — the Brownian motion and geometric Brownian motion — plateau early at a context length of about $10^{2}$, as shown in Figure 15. We attribute this early plateauing to the fact that the Brownian and geometric Brownian motions “wander out of distribution” at large $t$, while all other dynamical systems studied in this paper converge to stable distributions (i.e., the invariant measure).
A Markovian system (stochastic or deterministic) governed by a transition rule $P(x_{t+1}|x_{t})$ is said to have an invariant measure $\pi(x)$ if
[…]
If a system is initialized by $\pi(x)$ and evolves according to $P$, then the distribution of states at the next step will still follow $\pi(x)$. This property makes $\pi$ an invariant or stationary distribution for the system. It has been shown that the logistic map and Lorenz systems in the chaotic regime converge almost surely to their respective invariant measure, regardless of the initialization (Strogatz, 2015).
For discrete Markov chains governed by a transition matrix $p$, the stationary distribution is defined as a discrete probability mass function, denoted by $\vec{\pi}$, such that
$\vec{\pi}=p\vec{\pi},$ (13)
which is analogous to the continuous case described by Equation 12. By definition, any non-negative right eigenvector of $p$ with eigenvalue $\lambda=1$ is a stationary distribution of $p$. (Sethna, 2021) showed that a valid transition matrix has at least one stationary distribution.
On the other hand, neither the Brownian nor the geometric Brownian motion has invariant distributions [^3] For stochastic systems, the invariant measure is sometimes referred to as the stationary distribution. on unbounded domains (e.g., when $\mathcal{X}=\mathbb{R}$). This can be seen from the marginalized distribution $P(x_{t})$ at time $t$. For the Brownian motion defined in Equation 4, the marginalized distribution of $x_{t}$ at time $t$ is a normal distribution:
$P(x_{t})=\frac{1}{\sqrt{2\pi\sigma^{2}t}}\exp\left(-\frac{(x_{t}-\mu t)^{2}}{2\sigma^{2}t}\right),$ (14)
while for the geometric Brownian motion defined in Equation 5, the marginalized distribution of $x_{t}$ is a log-normal distribution (Crow & Shimizu, 1987):
$P(x_{t})=\frac{1}{x_{t}\sqrt{2\pi\sigma^{2}t}}\exp\left(-\frac{(\log(\frac{x_{t}}{x_{0}})-\mu t-\frac{\sigma^{2}}{2}t)^{2}}{2\sigma^{2}t}\right).$ (15)
Both Equations 14 and 15 are time-dependent and do not converge to a stationary distribution in the limit $t\rightarrow\infty$.
For the Brownian and geometric Brownian motions, the LLM might decide to only consider the most recent segment of time steps, and ignore the earlier data, which are in some sense “out of distribution”. This could explain the early plateauing of loss curves. Indeed, the classical neural scaling laws can be improved or broken if the scheduling of the training data shifts in distribution, as shown in (Sorscher et al., 2023). Different from (Sorscher et al., 2023; Lu et al., 2023), which alter the scheduling of data to achieve better learning curves that decrease faster with the size of training data, our experiments consider time series with pre-determined transition laws. We cannot tamper with the scheduling of our data to make it stationary without altering the underlying transition rules.
A.3 Temperature and variance
The temperature $T$ is a hyper-parameter that controls the variance of the softmax output layer. Although most LLMs are trained at $T=1$, it is common practice to tune the temperature in the interval $T\in[0.8,1.2]$ during inference. Then, one can opt for increased diversity (high $T$), or better coherence (low $T$) in the generated output. The temperature hyper-parameter affects the uncertainty, or variance, in the Hierarchy-PDF extracted from the LLM. Figures 17 and 17 show how different temperatures change the shape of the Hierarchy-PDF. In both cases, higher temperature leads to higher variance in the PDF.
We highlight the different refinement schemes used in these figures: for GBM, the PDF is refined to the last (third) digit near the mode, and left coarse elsewhere. This is because the true variance for GBM can span two orders of magnitude (see Figure 11), with most data points trapped in the low-variance region at small $X_{t}$. Hence, we require high precision to resolve these small variances in Figure 18. On the other hand, the noisy logistic map time series does not suffer from this issue, and thus we uniformly refine its PDF only up to the second digit.
Figure 16: Next state prediction for Geometric Brownian motion. Topmost: Input stochastic time series (black), and the state to be predicted (red). Rest: The Hierarchy-PDF prediction extracted from LLaMA-13b evaluated at different temperatures ranging from $T=0.3$ to $3$, along with ground truth PDF (red).
Figure 16: Next state prediction for Geometric Brownian motion. Topmost: Input stochastic time series (black), and the state to be predicted (red). Rest: The Hierarchy-PDF prediction extracted from LLaMA-13b evaluated at different temperatures ranging from $T=0.3$ to $3$, along with ground truth PDF (red).
Figure 17: Next state prediction for the noisy logistic map. Topmost: Input stochastic time series (black), and the state to be predicted (red). Rest: The Hierarchy-PDF prediction extracted from LLaMA-13b evaluated at different temperatures ranging from $T=0.3$ to $3$, along with ground truth PDF (red).
Figure 18: Most data points in GBM are trapped in low variance region with small $X_{t}$. The hierarchy-PDF must be very refined to resolve these minuscule variances.
While the loss curves in our paper are calculated at $T=1$, the predicted $\sigma$ shown in Figures 11 and 13 are extracted at $T=0.7$. We performed a grid search on the temperature ranging from $T=0.3$ to $T=3$ (see Figures 20 and 20), and observed that $T=0.7$ consistently results in better prediction quality of the variance.
Figure 19: GBM standard deviation $\sigma$ as a function of state value $X_{t}$, learned by the LLM, along with the ground truth. The LLM prediction is evaluated at temperatures ranging from $T=0.3$ to $1$.
LLM temperature = 0.3 Refer to caption LLM temperature = 0.7 Refer to caption LLM temperature = 1 Refer to caption
LLM temperature = 0.3 Refer to caption LLM temperature = 0.7 Refer to caption LLM temperature = 1 Refer to caption
Figure 19: GBM standard deviation $\sigma$ as a function of state value $X_{t}$, learned by the LLM, along with the ground truth. The LLM prediction is evaluated at temperatures ranging from $T=0.3$ to $1$.
Figure 20: Noisy logistic map standard deviation $\sigma$ as a function of state value $X_{t}$, learned by the LLM, along with the ground truth. The LLM prediction is evaluated at temperatures ranging from $T=0.3$ to $1$.
A.4 Non-linear pendulum
A non-linear pendulum is a classical physical system that extends the simple pendulum model by considering the effects of large angular displacements. Unlike the linear approximation, which assumes small oscillations and uses a simple harmonic oscillator equation, the non-linear pendulum model incorporates the full trigonometric term, which leads to complex non-linear phenomena such as amplitude-dependent frequency:
$\ddot{\theta}=-\lambda\dot{\theta}-\omega^{2}\sin(\theta).$ (16)
In the simulation shown in Figure 1, we use the initial condition $(\theta,\dot{\theta})=(0,1.05)$, the natural frequency $\omega=2$, and damping coefficient $\lambda=0.1$. The non-zero damping results in both amplitude and frequency drifting over time.
A.5 Markov chains with larger numbers of states and bigger LLM
Our experiments show that LLMs generally achieve lower in-context loss for Markov chains with fewer discrete states $n$, as shown in Figure 21. For both LLaMA-13b and LLaMA-70b, the in-context loss curves cease to decrease significantly for numbers of states $n\geq 9$.
Figure 21: In-context loss curves for LLaMA-13b (left) and LLaMA-70b (right) with respect to the number of states in the transition matrix.
A.6 Additional loss functions
KL-divergence. The KL-divergence between two PDFs, $P$ and $\tilde{P}$, is defined as
$D_{KL}(P,\tilde{P})=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{\tilde{P}(x)}\right).$ (17)
Although commonly used as the training loss for a variety of machine learning systems, this loss function may suffer from numerical instabilities as the learned transition function $\tilde{P}$ are often close to zero, as shown in Figure 5 and Figure 8, where the probability density is concentrated in small regions of the support.
Discretized Bhattacharyya distance for deterministic systems. For deterministic systems, the ground truth transition function is a delta function. Therefore, the Bhattacharyya distance between it and the Hierarchy-PDF prediction only references the finest bin associated with the true value $x_{\mathrm{true}}$.
$D_{B}(\delta(x-x_{\mathrm{true}}),\tilde{P})$ $=-\ln\left(\sum_{x}\sqrt{\delta(x-x_{\mathrm{true}})\tilde{p}(x)}\Delta x\right)=-\frac{1}{2}\ln(\tilde{p}(x_{\mathrm{true}}))-\ln\Delta x$ (18)
$=-\frac{1}{2}\ln(\tilde{p}(x_{\mathrm{true}}))+\text{constant}.$
As a result, the Bhattacharyya distance is reduced to an affine-transformed negative log-likelihood assigned to data by the LLM. Such local sensitivity on $\tilde{p}(x_{\mathrm{true}})$ explains the wild fluctuations seen in the Bhattacharyya loss in Section 4.1.
Higher moments and kurtosis. While the Bhattacharyya distance and SDM measure the agreement between the extracted transition rules $\tilde{P}$ and the ground truth distribution $P$, they do not explicitly characterize the type of the distribution (e.g., Gaussian or uniform). We employ the kurtosis as an additional measure to assess whether the LLM recovers the correct shape of $P$. The kurtosis of a distribution $P$ is defined as (Joanes & Gill, 1998)
$\text{Kurt}(P)=\frac{\mathbb{E}_{x\sim P}[(x-\mu_{P})^{4}]}{\mathbb{E}_{x\sim P}[(x-\mu_{P})^{2}]^{2}}=\frac{\mu_{4}}{(\sigma^{2})^{2}},$ (19)
where $\sigma^{2}$ and $\mu_{4}$ are the second and fourth central moments, which can be approximated using a hierarchical PDF as
$\sigma^{2}(P)=\sum_{x}p(x)(x-\mu_{p})^{2}\Delta x,\quad\mu_{4}(P)=\sum_{x}p(x)(x-\mu_{p})^{4}\Delta x.$ (20)
The kurtosis is equal to $3$ for a Gaussian distribution, and $\frac{9}{5}$ for bounded uniform distributions. Figure 22 shows the kurtosis of Brownian motion transition rules learned by LLM, which indeed converges to $3$ with increasing context length.
Figure 22: Kurtosis of Brownian motion transition rules with respect to the input length. Blue: kurtosis of LLM predicted PDF. Red: ground truth kurtosis, which is $3$ for all Gaussian distributions.
A.7 Hierarchy PDF
This section documents all three parts of the Hierarchy-PDF algorithm. We refer to the GitHub repository for further details.
Algorithm 2 Refine Each State in a Stochastic Sequence
  Input:
  - $S_{\text{traj}}$: A string representing a sampled stochastic trajectory whose states are separated by commas.
  - $L_{\text{PDF}}$: List of unrefined PDFs for each state.
[…]
  for each state and PDF in $S_{\text{traj}}$ and $L_{\text{PDF}}$ do
[…]
  end for
Algorithm 3 Detailed Hierarchy-PDF Recursive Refiner
  Input: Object multi_PDF representing unrefined PDF using bins of various widths
  Procedure: RecursiveRefiner(mainBranch, sequence, $D_{c}$, $D_{t}$, $KV_{\text{cache}}$)
     if $D_{c}=D_{t}$ then
        return {Terminate if target refinement depth is reached}
     end if
     if mainBranch is True then
        {Launch 9 recursive branches if the current sequence is refined}
        $L_{\text{new}}\leftarrow$ Form 9 new sequences by changing the last digits
        for each sequence in $L_{\text{new}}$ do
           RecursiveRefiner(False, $\texttt{sequence},D_{c},D_{t},KV_{\text{cache}}$)
        end for
     else
        {Collect refined logits}
        newLogits, newKVcache $\leftarrow$ NextTokenProbs($\texttt{sequence},KV_{\text{cache}}$)
        Refine multi_PDF using newLogits
     end if
     if $D_{c}+1<D_{t}$ then
        {Launch 10 more branches if $D_{t}$ not met}
        $L_{\text{new}}\leftarrow$ Form 10 new sequences by appending digits
        for each sequence in $L_{\text{new}}$ do
           RecursiveRefiner(False, $\texttt{sequence},D_{c}+1,D_{t},\texttt{newKVcache}$)
        end for
     end if
Algorithm 4 Extract Next Token Probabilities
  function NextTokenProbs(sequence, KVcache, model)
[…]
     Update KVcache
  return NextTokenLogit, KVcache